OpenStudy (anonymous):

***medal*** Let f(x) = 2x − 5 and g(x) = 7x − 3. What is (f(g(x))?

OpenStudy (anonymous):

I know the answer, would you like me to explain it?

OpenStudy (anonymous):

yes, i take a horal text tomorrow and i need to understand

OpenStudy (anonymous):

The answer is......something lol jk anyways I don't know but maybe ask @Teddyiswatshecallsme or @whatdoesthismean

OpenStudy (anonymous):

okay, tks

OpenStudy (anonymous):

@Teddyiswatshecallsme @whatdoesthismean

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

If you have f(g(x)), all you have to do is put in g(x) anywhere you see x in f(x).

OpenStudy (anonymous):

Your original equations are: f(x) = 2x − 5 and g(x) = 7x − 3 \[f(x)=2(7x-3)-5\] Now just solve for x.

OpenStudy (anonymous):

f(2)=2(7-3)-5

OpenStudy (anonymous):

its right?

hero (hero):

Actually, @adrianna45 \(f(x)\) does not equal \(2(7x - 3) - 5\)

OpenStudy (anonymous):

My mistake, \[f(g(x))=2(7x-3)-5\]

OpenStudy (anonymous):

so, i did wrong ;/